Listed below are the numbers of home runs in major league baseball for each year beginning with 1990.

3317 3383 3038 4030 3306 4081 4962 4640 5064 5528

Find the median

Find the variance
Answer #1

Solution:

x x2
3317 11002489
3383 11444689
3038 9229444
4030 16240900
3306 10929636
4081 16654561
4962 24621444
4640 21529600
5064 25644096
5528 30558784
∑x=41349 ∑x2=177855643

a ) Median :
Observations in the ascending order are :
3038,3306,3317,3383,4030,4081,4640,4962,5064,5528

Here, n=10 is even.

M=Value of(n/2)thobservation+Value of(n2+1)thobservation2

=Value of(10/2)thobservation+Value of(10/2+1)thobservation2

=Value of5thobservation+Value of6thobservation2

=4030+40812

=4055.5

b ) Sample Variance S2=∑x2-(∑x)2nn-1

=177855643-(41349)210/9

=177855643-170973980.1/9

=6881662.9/9

=764629.2111